# Multimodal Biomarkers for Schizophrenia: Towards Individual Symptom Severity Estimation

### v1

@@ 1 Introduction
Schizophrenia is a complex and heterogeneous mental health disorder that affects approximately 24 million people worldwide [1]. Despite its significant impact, many individuals with schizophrenia and other mental health conditions remain undiagnosed or experience delays in receiving appropriate treatment. Limited access to mental healthcare, particularly in underserved and rural areas, exacerbates this issue, as mental health specialists and resources are often scarce [2]. Even when care is accessible, the diverse and overlapping symptoms of schizophrenia pose a diagnostic challenge, frequently leading to misdiagnosis or prolonged delays in treatment. Traditional diagnostic approaches rely heavily on subjective clinical assessments, which can vary between practitioners and lack consistency. These challenges underscore the urgent need for accessible, scalable, and objective measures that can aid in symptom assessment by identifying reliable biomarkers and clinical subtypes that can capture pathological homology missed by conventional diagnostic assessments. Advancements in this area would support early and accurate identification and characterization of schizophrenia symptom profiles, ultimately yielding meaningful neurobiological and behavioral targets for treatments, etiological investigations and patient outcomes.
One promising avenue for objective mental health assessment lies in multimodal analysis, integrating speech, video,and text-based inputs[3, 4]. Recent research has demonstrated that speech patterns, facial expressions, and linguistic features extracted from transcripts can serve as valuable, non-invasive indicators of mental health conditions, particularly in disorders such as major depressive disorder [5], post-traumatic stress disorder [6], and schizophrenia [7]. Accordingly, these multimodal objective signals can be used in order to identify clinical subtypes that converge on symptom domains as assessed by more subjective clinical interviews. Speech-based methods analyze vocal patterns, acoustic features, and articulatory characteristics, while video analysis can capture facial expressions and gaze behavior indicative of underlying symptoms. Additionally, text from transcripts of natural speech or clinical interviews can provide insights into linguistic coherence, thought patterns, and semantic anomalies associated with schizophrenia. Integrating speech, video, and text-based modalities could improve the robustness and accuracy of symptom detection, allowing clinicians to tailor treatment strategies based on specific symptom severity.
@@
One critical consideration in developing such systems is whether to use multiple models for different symptoms or a single unified model capable of estimating all symptoms simultaneously. A unified model offers several advantages. First, schizophrenia symptoms often co-occur and influence one another, meaning that a model trained to recognize all symptoms collectively may better capture these interdependencies. Second, training and deploying separate models for each symptom would be computationally expensive and inefficient, requiring significant resources for both development and real-time inference. Third, a single model ensures greater consistency in symptom estimation, reducing variability that might arise when combining predictions from multiple independent models. Lastly, integrating all symptoms into a unified framework allows for a more holistic and clinically meaningful assessment, aligning more closely with real-world diagnostic practices where clinicians evaluate symptoms in conjunction rather than in isolation.
@@ 2 Dataset
The dataset used in this study is a dataset that was collected as a part of an interdisciplinary mental health assessment project jointly carried out by the University of Maryland School of Medicine and the University of Maryland - College Park [9]. The dataset contains audio and video recordings of subjects who were interviewed during their in-person clinic visits. The dataset contains data from a total of 140 sessions produced by 39 different subjects. Before their interview sessions they were also evaluated based on mental health assessment questionnaires like Hamilton Depression rating scale (HAMD)[10] and Brief Psychiatric Rating Scale (BPRS) [11] to quantify their symptom severity before their sessions. In this study we are focusing our experiments based on the 18-item BPRS scores as they look into the psychiatric symptoms presented by the subjects which are good indicators for schizophrenia. In the 18-item BPRS scale all of these symptoms are individually rated on a scale of 1-7 that goes up with increasing severity of the particular symptom. For example, a score of 1 stands for symptom not present, a score of 4 meaning moderate severity, and a score of 7 stands for extremely severe symptom. Table.1 gives a summary of the 18 symptoms and the frequency of the symptom scores across all the sessions in the dataset.
@@ 3.2 Speech-based feature extraction
For speech based features, previous works on mental health assessment have shown that different features like pre-trained self-supervised acoustic features, articulatory coordination features and concise articulatory representations have shown promising performance in speech-based assessment across various tasks. So in this work, we extracted all these features and tested them out on different unimodal models. For pre-trained self-supervised acoustic features we extracted Wav2Vec2.0 [12] and WavLM [13] features. For articulatory features, we extracted vocal tract variables (TVs) from an acoustic to articulatory speech inversion system [14] along with source articulatory features from an Aperiodicity, Periodicity, Pitch detector[15]. And then converted them into Full Vocal Tract coordination (FVTC) features using a time-delay channel correlation mechanism [16]. To obtain Concise Articulatory (C-Art) representations we used a vector Quantized Variational Auto Encoder (VQ-VAE)-based representation learning model on FVTC features [8]. For session-level representations, the extracted features from all 40-second segments of the session were stacked together.
Figure 1: Multimodal model architecture with Wav2vec2.0, C-Art represntations, BERT embeddings, and FAU-Coordination features.
@@ 3.3 Video-based feature extraction
For video-based features, we extracted Facial Action Units (FAU) [17] from Openface toolbox [18], the FAUs capture the movement of various facial muscles and facial landmarks. After extracting the FAUs, the FAU-based Coordination (FAUC) features were calculated similar to the way FVTC features were calculated from TVs using the same time-delay channel correlation mechanism [16]. To produce session-level representations, the extracted FAU-based coordination features from all 40-second segments of the session were stacked together.
@@ 3.5 Unimodal models
Unimodal models were independently trained and evaluated for speech, video, and text inputs. For speech-based analysis, multiple models were developed using pre-trained self-supervised acoustic features, FVTC features, and C-Art representations. Given previous findings [8] that fusing pre-trained acoustic features with C-Art representations improves performance on mental health assessment tasks, we also incorporated the Acoustic-Articulatory (AA)-Fusion approach in our speech-based unimodal model training and evaluation. For video and text inputs, unimodal models were trained using facial action unit (FAU)-based coordination features and BERT embeddings respectively. All the unimodal model architectures were designed with a Convolution Neural Network (CNN) backbone followed by a set of linear layers for multi-label multi-classs predictions. The CNN style architecture was chosen because the stacked unimodal representations produced 2-dimensional inputs which can be effectively analyzed using the filters in CNN to capture spatial relationships and patterns among them.

### v2

@@ 1 Introduction
Schizophrenia is a complex and heterogeneous mental health disorder that affects approximately 24 million people worldwide [1]. Despite its significant impact, many individuals with schizophrenia and other mental health conditions remain undiagnosed or experience delays in receiving appropriate treatment. Limited access to mental healthcare, particularly in under-served and rural areas, exacerbates this issue, as mental health specialists and resources are often scarce [2]. Even when care is accessible, the diverse and overlapping symptoms of schizophrenia pose a diagnostic challenge, frequently leading to misdiagnosis or prolonged delays in treatment. Traditional diagnostic approaches rely heavily on subjective clinical assessments, which can vary between practitioners and lack consistency. These challenges underscore the urgent need for accessible, scalable, and objective measures that can aid in symptom assessment by identifying reliable biomarkers and clinical subtypes that can capture pathological homology missed by conventional diagnostic assessments. Advancements in this area would support early and accurate identification and characterization of schizophrenia symptom profiles, ultimately yielding meaningful neurobiological and behavioral targets for treatments, etiological investigations and patient outcomes.
One promising avenue for objective mental health assessment lies in multimodal analysis, integrating speech, video,and text-based inputs[3, 4, 5]. Recent research has demonstrated that speech patterns, facial expressions, and linguistic features extracted from transcripts can serve as valuable, non-invasive indicators of mental health conditions, particularly in disorders such as major depressive disorder [6], post-traumatic stress disorder [7], and schizophrenia [8]. Accordingly, these multimodal objective signals can be used in order to identify clinical subtypes that converge on symptom domains as assessed by more subjective clinical interviews. Speech-based methods analyze vocal patterns, acoustic features, and articulatory characteristics, while video analysis can capture facial expressions and gaze behavior indicative of underlying symptoms. Additionally, text from transcripts of natural speech or clinical interviews can provide insights into linguistic coherence, thought patterns, and semantic anomalies associated with schizophrenia. Integrating speech, video, and text-based modalities could improve the robustness and accuracy of symptom detection, allowing clinicians to tailor treatment strategies based on specific symptom severity.
@@
One critical consideration in developing such systems is whether to use multiple models for different symptoms or a single unified model capable of estimating all symptoms simultaneously. A unified model offers several advantages. First, schizophrenia symptoms often co-occur and influence one another, meaning that a model trained to recognize all symptoms collectively may better capture these inter-dependencies. Second, training and deploying separate models for each symptom would be computationally expensive and inefficient, requiring significant resources for both development and real-time inference. Third, a single model ensures greater consistency in symptom estimation, reducing variability that might arise when combining predictions from multiple independent models. Lastly, integrating all symptoms into a unified framework allows for a more holistic and clinically meaningful assessment, aligning more closely with real-world diagnostic practices where clinicians evaluate symptoms in conjunction rather than in isolation.
@@ 2 Dataset
The dataset used in this study is a dataset that was collected as a part of an interdisciplinary mental health assessment project jointly carried out by the University of Maryland School of Medicine and the University of Maryland - College Park [10]. The dataset contains English audio and video recordings of subjects who were interviewed during their in-person clinic visits. The dataset contains data from a total of 140 sessions produced by 39 different subjects. Before their interview sessions they were also evaluated based on mental health assessment questionnaires like Hamilton Depression rating scale (HAMD)[11] and Brief Psychiatric Rating Scale (BPRS) [12] to quantify their symptom severity before their sessions. In this study we are focusing our experiments based on the 18-item BPRS scores as they look into the psychiatric symptoms presented by the subjects which are good indicators for schizophrenia. In the 18-item BPRS scale all of these symptoms are individually rated on a scale of 1-7 that goes up with increasing severity of the particular symptom. For example, a score of 1 stands for symptom not present, a score of 4 meaning moderate severity, and a score of 7 stands for extremely severe symptom. Table.1 gives a summary of the 18 symptoms and the frequency of the symptom scores across all the sessions in the dataset.
@@ 3.2 Speech-based feature extraction
For speech based features, previous works on mental health assessment have shown that different features like pre-trained self-supervised acoustic features, articulatory coordination features and concise articulatory representations have shown promising performance in speech-based assessment across various tasks. So in this work, we extracted all these features and tested them out on different unimodal models. For pre-trained self-supervised acoustic features we extracted Wav2Vec2.0 [13] and WavLM [14] features. For articulatory features, we extracted vocal tract variables (TVs) from an acoustic to articulatory speech inversion system [15] along with source articulatory features from an Aperiodicity, Periodicity, Pitch detector [16]. These TVs were then converted into Full Vocal Tract coordination (FVTC) features using a time-delay channel correlation mechanism [17]. Previous studies has been shown that changes in mental health conditions affect the coordination between articulatory gestures and these coordination changes are captured using these correlation features [18]. In addition, since the time-delay correlation matrices are sparse and contain redundant information, we extracted a Concise Articulatory (C-Art) representations using a vector Quantized Variational Auto Encoder (VQ-VAE)-based representation learning model on the FVTC features [9]. For session-level representations, the extracted features from all 40-second segments of the session were stacked together and padded to match the duration of the longest session in the dataset.
Figure 1: Multimodal model architecture with Wav2vec2.0, C-Art representations, BERT embeddings, and FAU-Coordination features.
@@ 3.3 Video-based feature extraction
For video-based features, we extracted Facial Action Units (FAU) [19] from Openface toolbox [20], the FAUs capture the movement of various facial muscles and facial landmarks. After extracting the FAUs, the FAU-based Coordination (FAUC) features were calculated similar to the way FVTC features were calculated from TVs using the same time-delay channel correlation mechanism [17]. To produce session-level representations, the extracted FAU-based coordination features from all 40-second segments of the session were stacked together and padded to match the duration of the longest session.
@@ 3.5 Unimodal models
Unimodal models were independently trained and evaluated for speech, video, and text inputs. For speech-based analysis, multiple models were developed using pre-trained self-supervised acoustic features, FVTC features, and C-Art representations. Given previous findings [9] that fusing pre-trained self-supervised acoustic features with C-Art representations for an Acoustic-Articulatory (AA)-Fusion approach improves performance on schizophrenia assessment task, we also trained and tested a similar AA-Fusion based speech model for this individual symptom severity estimation. For video and text inputs, unimodal models were trained using facial action unit (FAU)-based coordination features and BERT embeddings respectively. All the unimodal model architectures were designed with a Convolution Neural Network (CNN) backbone followed by a set of linear layers for multi-label multi-class predictions. The CNN style architecture was chosen because the stacked unimodal representations produced 2-dimensional inputs which can be effectively analyzed using the filters in CNN to capture spatial relationships and patterns among them.
@@ 6 Conclusion & Future work
7 Acknowledgment
This work was supported by the National Science Foundation grant numbered 2124270.